Hi all see this 2 Questions plz
1- A die was rolled 720 times , what is the probability that the number 4 will be exactly 125 times ?
2- What is the probability of getting at most three heads in 8 flips of balanced coin ?

Answer by stanbon(57250)

1- A die was rolled 720 times , what is the probability that the number 4 will be exactly 125 times ?
Prob(X=125)= [720C125](1/6)^125(5/6)^595=0.03466...
Comment: If you are using a TI-83 the problem will cause an
"overflow" when you try to calculate 720C125. However, if
you use binompdf(720,1/6,125) you will get 0.03466...
-----------------

2- What is the probability of getting at most three heads in 8 flips of balanced coin ?
P(0 Comment: You could also get this answer by calculating P(X=0)
+P(X=1)+P(X=2)+P(X=3)
--------
Hope this helps.
Cheers,
Stan H.